Calculate the following line integral:
∫ C xy^3dx, where C is the unit circle x^2+y^2=1, oriented counterclockwise. Note: C is at the bottom right of the integral.
1
Expert's answer
2019-11-12T10:11:06-0500

Counterclockwise parametrization of "C" is "(\cos t,\sin t)" , "t\in[0,2\pi]" , so

"\int\limits_C xy^3dx=\int\limits_0^{2\pi} \cos t\sin^3 td(\cos t)="

"=-\int\limits_0^{2\pi} \cos t\sin^4 tdt=-\int\limits_0^{2\pi}\sin^4 td(\sin t)=-\frac{1}{5}\sin^5 t\bigl|_0^{2\pi}="

"=-\frac{1}{5}\sin^5 2\pi+\frac{1}{5}\sin^5 0=0"

Answer: